OpenStudy (anonymous):

how many 5-letter arrangements of the letters R E G I O N A L consisting of 3 consonants and 2 vowels can be formed if no letter is repeated?

OpenStudy (anonymous):

How many consonants and vowels are there?

OpenStudy (anonymous):

4 consonants and 4 vowels

OpenStudy (anonymous):

In how many ways you can arrange 2 vowels out of 4? In how many ways you can arrange 3 consonant out of 4?

OpenStudy (anonymous):

so 4C2 * 4C3?

OpenStudy (anonymous):

well first you select first 2 vowels from 4 so thats: 4!/2!2! then you select 3 consonants from 4: 4!/2!2!+4!/3!1! and then we arrange them so thats just 5! because for the first letter word, we have 5 choices, for 2nd, since rep. not allowed 4 choices, for 3rd, 3 choices, 4th, 2 choices, 5th one choice, 1. So we'll have 4C2+4C3+5!

OpenStudy (anonymous):

I don't get how you got the 5 factorial and why are you adding these?

OpenStudy (anonymous):

288 is the answer?

OpenStudy (anonymous):

I don't know I'm still not sure about this :(

OpenStudy (anonymous):

I think it is (number of vowel combinations)* (number of consonant combinations)*(number of 5 letter arrangements), so 4C2*4C3*5! should be correct

OpenStudy (anonymous):

I don't get it, am I supposed to add or subtract?

OpenStudy (anonymous):

you shoud multiply, combinations work that way: for example (5 combinations)*(3 combinations)=(c1+c2+c+c4+c5)*(comb1+comb2+comb3)= 15 diferent options

OpenStudy (anonymous):

thank you!